# Leveraging Structural Knowledge in Diffusion Models for Source Localization in Data-Limited Graph Scenarios

### v2

@@ 4.2. Structure-prior biased Denoising Process
In most propagation scenarios, source nodes constitute a small minority of total graph nodes (Cheng et al., 2024). With limited training data, this severe class imbalance often leads models to degenerate towards predicting all nodes as non-sources, as the model overfits to the dominant non-source class. To address this, inspired by Han et al. (2022), we propose an structure-prior biased denoising scheme that set the prior of the denoising process to structure-based source estimations rather than random noise (illustrated in Figure 2). This design leverages two key insights: (1) structure-based estimations provide reliable initial source candidates by capturing inherent source structural properties like centrality, making them more informative than random noise, and (2) starting the denoising process from these estimations creates a natural bias in the trajectory space, encouraging the model to explore regions with higher likelihood of containing true sources. By incorporating these topology-aware priors into the initialization and denoising process, our approach effectively prevents model collapse while reducing the data requirements for learning accurate source patterns.
@@
(6) $p(X_{t}|X_{0},Y,\mathcal{G}))=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}X_{0}+(1-\sqrt{\bar{\alpha}_{t}})X_{est}(Y,\mathcal{G}),(1-\bar{\alpha}_{t})I),$
@@
(8) $\tilde{\mu}(X_{t},X_{0},X_{est}(Y,\mathcal{G})):=$ $\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}X_{0}+\frac{(1-\bar{\alpha}_{t-1})\sqrt{\alpha_{t}}}{1-\bar{\alpha}_{t}}X_{t}+$
@@
$\tilde{\beta}_{t}:=$ $\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}.$
The denoising network $f_{\theta}$ is set to estimate the ground truth source $X$(i.e. $X_{0}$ in Equation (8)), which we empirically find more effective. The denoise network outputs the estimated source vector $\tilde{X}_{0}:=f_{\theta}(X_{t},X_{est},Y,\mathcal{G},t)$ to calculate the posterior for step-by-step denoising. The denoising network $f_{\theta}$ can be trained by the simple L2 loss function (Ho et al., 2020):
(9) $L(\theta)=\mathrm{E}_{X_{0}\sim p(X_{0}|\cdot),t,\epsilon}||X_{0}-f_{\theta}(X_{t},t,\cdot)||^{2}_{2},$
@@ 4.3.1. Denoising Network Architecture
Encoding the noisy input and soft labels. The pre-estimated $X_{est}$ is forwarded through a multi-layer GNN to capture the hidden message with graph structural information. Subsequently, it is added to the noisy input $X_{t}$ and passed through a linear layer. The final input for the GNN encoder is $Z_{e}=\mathrm{Linear}(\mathrm{GNN}(X_{est})\oplus X_{t})\oplus Emb(t),$ where for the denoising step t, we use the classical sinusoidal embedding (Vaswani et al., 2017). The $\oplus$ indicates element-wise sum. $Z_{e}$ is then passed through a GCN-based encoder and is smoothed through a softmax function $\sigma$ and layer normalization:
(10) $Z_{d}=\mathrm{LN}(\sigma(\mathrm{GNN}(Z_{e}))).$
@@
Conditioning. Shown at the left part of the figure, a GCN-based module learns the encoding carrying the source prominence and centrality from the infection state input $Y$, which will be elaborated on in the next section.
Decoder. $Z_{d}$ and encoded condition $h_{out}$ are decoded through a GCN-based module, resulting in the estimation for the uncorrupted sample $X_{0}$ (i.e. $X$):
(11) $\tilde{X_{0}}=\mathrm{GNN}(Z_{d},h_{out}).$
@@ 4.3.2. GNN-parameterized Label Propagation(GNN-LP)
Our conditioning module takes the observed infection states as input, which contain crucial coupling information between spreading dynamics and network topology. To effectively extract this information, we introduce the GNN-LP module. We first employ label propagation (Wang et al., 2017), which first mark infected nodes as 1 and uninfected as -1 in the observation state $Y$, resulting in $Y^{*}$. The propagation follows the update rule: the label of a node in the next step is a combination of its original label and the sum of normalized labels from its neighbors. We can rewrite this iteration as:
@@
where we add non-linear transformations $h(\cdot)$ and $\sigma(\cdot)$ to enhance the expressiveness of the propagation process. As propagation continues, each node’s label value changes according to the graph structure. These changes reflect the node’s importance (centrality) and spreading influence in the network. By observing these changes, we can understand the relationship between propagation patterns and network structure. We then enhance this structural diffusion with GNNs, which learn topology-specific message passing rules to capture how different local structures influence propagation patterns. We can notice that the structure of the above equation exactly matches the form of the general Graph Neural Network (GNN) (Gilmer et al., 2017), and the parameterization can be achieved by using a residual block combined with a graph convolutional network(GCN, (Kipf and Welling, 2016)):
@@
2:// Phase 1: Learn from Diverse Propagation Patterns
3:Generate synthetic $(X^{*},Y)$ pairs using $\mathcal{M}s$ on $\mathcal{G}$
4:Train diffusion model using Eq.9 with $(X^{*},Y,\mathcal{G})$
5:// Phase 2: Adapt to Real Scenarios
6:for each $(X^{*},Y)$ in $\mathcal{D}_{real}$ do
7:     Fine-tune model using Eq.9 with $(X^{*},Y,\mathcal{G})$
8:end for
9:// Inference on Real New Cases
10:Input: Observation $Y_{new}$ on $\mathcal{G}$
11:Compute structure-based prior $X_{est}$ from $Y_{new}$ using Eq.1
12:Sample initial noise $X_{n}\sim\mathcal{N}(0,I)$
4.4. Pretrain using Simulation data from Established Models
The proposed structure-prior informed diffusion framework effectively extracts topology-aware features that remain largely stable across different propagation dynamics. These designs enable the framework to learn diverse knowledge from different propagation patterns while maintaining stable, pattern-invariant feature extraction capabilities centered on network structure. With this ability to capture generalizable source-topology relationships, we can effectively utilize synthetic data from established models (e.g. IC, LT, SIS) for pretraining, followed by efficient adaptation to real-world scenarios through few-shot learning, thus providing a practical solution for source localization in settings where real-world propagation training data is limited. We provide the illustration of the above process in Algorithm 1, and in Section 5.3, we evaluate SIDSL’s ability to bridge simulation and reality by comparing two approaches: one using pretrained models with few-shot learning, and another trained exclusively on real-world data.
@@ 5.1. Experiment Settings
Datasets. Following Ling et al. (2022) and  Huang et al. (2023b), we use four real-world propagation data to evaluate SIDSL. The real-world datasets include Digg, Twitter, Android and Christianity, in which the real propagation cascades are available. For each cascade in all sets, we designate the infected nodes at the first 10% of the propagation time as source nodes and take the network’s infection status at 30% of the propagation time as observation input. In the context of real-world applications, we often can only collect sufficient data for analysis after some time has elapsed since the occurrence of the event. Therefore, attempting to predict what initially happened in the process when we have observed enough propagation patterns at a certain degree of infection time is very much in line with the needs of real-world operations. Please refer to the Appendix D for specific details of the datasets.
Implementation Details. For each dataset, the ratio of training, validation, and testing portion is 6:1:1. For the diffusion framework of SIDSL, we use $T=500$ maximum diffusion timestep and linear schedule for noise scheduling. In the denoising network, we leverage a 2-layer graph convolutional network (GCN) to forward the LPSI estimation $X_{est}$ condition. The GNN encoder and decoder comprise 3-layer GCNs with a hidden dimension of 128. The residual GNN of the conditioning module is a 2-layer GCN, with a hidden dimension of 8. The learning rate is searched from 0.01, 0.005, 0.001, and the maximum number of training epochs is set to 500 for all datasets. We train our model using Adam optimizer and a learning rate scheduler with a linear decay. Our model is trained on a single NVIDIA GeForce RTX 2080 Ti. The code implementation can be found at https://anonymous.4open.science/r/ASLDiff-4FE0.
Baselines. Following previous works (Ling et al., 2022; Yan et al., 2024), we selected two representative heuristic methods, i.e., Netsleuth (Prakash et al., 2012) and LPSI (Wang et al., 2017), and deep learning methods, i.e., GCNSI (Dong et al., 2019), TGASI (Hou et al., 2023) and recent generative deep learning methods SLVAE (Ling et al., 2022), DDMSL (Yan et al., 2024). These baselines are all state-of-the-art (SOTA) multi-source localization methods in their domains. Please refer to the Appendix E for specific information of baselines and our method.
Metrics. Following previous works (Wang et al., 2022), we adopt three metrics: 1) F1-score (F1): The harmonic mean of recall and precision, emphasizing the balance between precision and recall; 2) Recall (RE): The proportion of positive cases (source nodes) that are correctly identified, focusing on the model’s ability to detect all relevant instances; 3) Precision (PR): The proportion of actual positive cases among the samples judged as positive, highlighting the model’s ability to avoid false positives. We do not use accuracy since in highly imbalanced datasets with rare positive cases, accuracy is a misleading metric as it can achieve deceptively high values by simply predicting the majority (negative) class, thus failing to capture the model’s true discriminative capability for the minority class (source) of interest.
@@ 5.2. Overall Performance on Real-world Datasets
To evaluate the real-world performance of our proposed method against baselines, we conduct direct training and testing on four real-world datasets for all methods. The experimental results are presented in Table 1. Our proposed SIDSL outperforms all baselines across nearly all metrics on all datasets. Specifically, SIDSL achieves F1 scores that exceed the second-best baseline by 7.5%, 11.0%, 11.8%, and 13.3% on Digg, Twitter, Android, Christianity, respectively, which demonstrates SIDSL’s superior ability to predict source nodes despite their sparsity.
Further, we have the following findings. First, compared to other generative methods like DDMSL and SLVAE which rely purely on data-driven approaches, SIDSL demonstrates superior performance, validating the effectiveness of incorporating structural prior information on top of purely data-driven distribution learning. Second, compared to rule-based and pure learning-based methods, generative methods (SIDSL, DDMSL and SLVAE) show relatively strong performance, highlighting the advantages of effective source distribution modeling for source localization accuracy. LPSI and other deep learning-based methods perform comparably well, as they successfully capture structural characteristics of source nodes through either rules or learning. Netsleuth demonstrates the weakest performance, likely due to its design for SI propagation patterns (Keeling and Eames, 2005; Prakash et al., 2012), which limits its generalization to more complex patterns and larger network scales in real-world data. Finally, comparing the performance across different datasets, the most substantial improvement of SIDSL is observed on the Christianity dataset (13.3% in F1), likely because its relatively smaller network scale (in terms of nodes and edges) makes topological features more easily capturable by all methods, thereby amplifying the benefits of structural prior knowledge injection in SIDSL.
@@ 5.3. Performance in Low Data Regime
To assess how deep learning methods leverage simulation data pretraining for few-shot and zero-shot learning in real-world source localization, we generate pretraining data using standard IC and LT propagation models (1:1 ratio) across four networks. These models effectively capture social network dynamics by representing real-world interaction randomness, cumulative influence, and topology-based spread patterns. For few-shot learning (P), we fine-tune the pretrained models with limited real data (10%). We also compare the performance of all the methods trained on equivalent data volumes but without pretraining (NP). For zero-shot learning, all the methods are directly tested after being pretrained.
@@
Few-shot analysis. The result of few-shot learning is shown in Tabel 2. Rule-based methods (Netsleuth and LPSI) are not shown here because they are training-free. SIDSL demonstrates superior performance compared to all baseline methods across all datasets, both with and without pretraining, achieving improvements of 24$\sim$365across all metrics. Further, we have the following two findings. First, SIDSL demonstrates enhanced performance advantages over baselines when leveraging pretraining, as evidenced by the improvement percentages across datasets. Among the improvements ($\Delta$) across different datasets and metrics, 8 out of 12 cases show larger gains with P compared to NP. This consistent pattern of larger improvements with pretraining demonstrates SIDSL’s superior ability to effectively leverage information from simulated data to real-world scenarios in limited-data settings. Second, the baseline methods exhibit varying capabilities in utilizing pretraining. While the non-generative method TGASI shows consistent improvements with pretraining across datasets, generative approaches like SLVAE and DDMSL demonstrate limited and unstable gains (e.g., SLVAE’s F1 only increases from 0.009 to 0.036 while DDMSL’s F1 only increases from 0.010 to 0.109 on Android). This contrast highlights the generative models’ sensitivity to distribution shifts between simulated and real-world data, whereas our method achieves robust and significant improvements through effective incorporation of pattern-invariant structural prior information.
@@
Zero-shot analysis. The results of zero-shot learning are shown in Table 3. SIDSL significantly outperforms all baseline methods across all datasets, achieving improvements of 85%3̃64% in F1 scores. The performance gap is particularly notable on Android and Christianity datasets, where SIDSL achieves over 0.9 recall while maintaining reasonable precision. In contrast, baseline methods show limited zero-shot transfer capability with F1 scores mostly below 0.3. Compared to rule-based methods in Table 1, zero-shot SIDSL still outperforms them in most metrics. These findings demonstrates SIDSL’s superior ability to leverage knowledge from simulation data without fine-tuning.
@@
Figure 4. Few-shot performance of SIDSL with different pretraining dataset compositions (IC+LT, IC, LT), followed by finetuning on real-world propagation data. ”NP” denotes ”non-pretrained”.

### v3

@@ 4.2. Structure-prior Biased Denoising Process
Source nodes constitute a small minority of total graph nodes (Cheng et al., 2024). With limited training data, this severe class imbalance leads models to degenerate towards predicting all nodes as non-sources. Following Han et al. (2022), we propose a structure-prior biased denoising scheme that sets the prior to structure-based source estimations rather than random noise (Figure 2). This leverages two insights: (1) structure-based estimations capture inherent source properties like centrality, providing more informative priors than random noise, and (2) initializing denoising from these estimations biases exploration toward regions with higher source likelihood. Formally, this approach decomposes the posterior as $p(X_{0}|Y,\mathcal{G})\propto p_{{prior}}(X_{{est}}|Y,\mathcal{G})\cdot p(X_{0}|X_{{est}})$, where the structural prior $p_{{prior}}$ anchors learning in high-probability source regions, enabling the denoiser $p(X_{0}|X_{{est}})$ to focus on lower-variance refinements that are learnable with scarce data. This modified forward process shifts the diffusion trajectory from converging to noise toward structure-informed regions, while preserving generative correctness as our sampling process matches the modified forward posterior’s functional form. These topology-aware priors prevent model collapse while reducing data requirements for accurate source learning.
@@
(6) $p(X_{t}|X_{0},Y,\mathcal{G})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}X_{0}+(1-\sqrt{\bar{\alpha}_{t}})X_{est}(Y,\mathcal{G}),(1-\bar{\alpha}_{t})I),$
@@
(8) $\tilde{\mu}(X_{t},X_{0},X_{est}(Y,\mathcal{G}))$= $\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}X_{0}+\frac{(1-\bar{\alpha}_{t-1})\sqrt{\alpha_{t}}}{1-\bar{\alpha}_{t}}X_{t}+$
@@
$\tilde{\beta}_{t}$= $\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}.$
This formulation requires that the denoiser $f_{\theta}$ learns to predict $X_{0}$ given the modified trajectory, ensuring the sampling posterior $q_{\theta}(X_{t-1}|X_{t},X_{est})$ matches the ground-truth form $p(X_{t-1}|X_{t},X_{0},X_{est})$ with $X_{0}$ replaced by the denoiser’s prediction. The denoising network outputs the estimated source vector $\tilde{X}_{0}:=f_{\theta}(X_{t},X_{est},Y,\mathcal{G},t)$ to calculate the posterior for step-by-step denoising. The denoising network $f_{\theta}$ can be trained by the simple L2 loss function (Ho et al., 2020):
[…]
@@ 4.3.1. Denoising Network Architecture
Encoding the noisy input and soft labels. The pre-estimated $X_{est}$ is forwarded through a multi-layer GNN to capture the hidden message with graph structural information. Subsequently, it is added to the noisy input $X_{t}$ and passed through a linear layer. The final input for the GNN encoder is $Z_{e}=\mathrm{Linear}(\mathrm{GNN}(X_{est})\oplus X_{t})\oplus Emb(t),$ where for the denoising step t, we use the classical sinusoidal embedding (Vaswani et al., 2017). The $\oplus$ indicates element-wise sum. $Z_{e}$ is then passed through a GCN-based encoder and is smoothed through a softmax function $\sigma$ and layer normalization: $Z_{d}=\mathrm{LN}(\sigma(\mathrm{GNN}(Z_{e}))).$
@@
Conditioning. Shown on the left part of the figure, a GCN-based module learns the encoding carrying the source prominence and centrality from the infection state input $Y$, which will be elaborated on in the next section.
Decoder. $Z_{d}$ and encoded condition $h_{out}$ are decoded through a GCN-based module, resulting in the estimation for the uncorrupted sample $X_{0}$ (i.e. $X$): $\tilde{X_{0}}=\mathrm{GNN}(Z_{d},h_{out}).$
@@ 4.3.2. GNN-parameterized Label Propagation(GNN-LP)
Our conditioning module takes the observed infection states as input, which contain crucial coupling information between spreading dynamics and network topology. To effectively extract this information, we introduce the GNN-LP module. We first employ label propagation (Wang et al., 2017), which first marks infected nodes as 1 and uninfected as -1 in the observation state $Y$, resulting in $Y^{*}$. The propagation follows the update rule: the label of a node in the next step is a combination of its original label and the sum of normalized labels from its neighbors. We can rewrite this iteration as:
@@
where we add non-linear transformations $h(\cdot)$ and $\sigma(\cdot)$ to enhance the expressiveness of the propagation process. During propagation, node labels evolve with the graph structure, reflecting node importance and the relationship between propagation patterns and network structure. We enhance this with GNNs, which learn topology-specific message passing rules to capture how local structures influence propagation. This method’s formulation matches the general GNN (Gilmer et al., 2017) framework and is parameterized by a GCN (Kipf and Welling, 2016) combined with a residual block:
@@
2:// Phase 1: Learn from Diverse Propagation Patterns
3:Generate synthetic $(X^{*},Y)$ pairs using $\mathcal{M}s$ on $\mathcal{G}$
4:Train diffusion model using Eq.9 with $(X^{*},Y,\mathcal{G})$
5:// Phase 2: Adapt to Real Scenarios
6:for each $(X^{*},Y)$ in $\mathcal{D}_{real}$ do
7:  Fine-tune model using Eq.9 with $(X^{*},Y,\mathcal{G})$
8:end for
9:// Inference on Real New Cases
10:Input: Observation $Y_{new}$ on $\mathcal{G}$
11:Compute structure-based prior $X_{est}$ from $Y_{new}$ using Eq.1
12:Sample initial noise $X_{n}\sim\mathcal{N}(0,I)$
4.4. Pretrain Using Simulation Data
The proposed structure-prior informed diffusion framework effectively extracts topology-aware features that remain stable across different propagation dynamics. This enables pretraining on synthetic data from established models (IC, LT, SIS) with efficient few-shot adaptation to real scenarios, addressing data-limited source localization. We illustrate this process in Algorithm 1 and evaluate SIDSL’s simulation-to-reality transfer in Section 5.3 by testing two approaches: one using pretrained models with few-shot learning, and another trained exclusively on real-world data.
@@ 5.1. Experiment Settings
Datasets. Datasets. Following Ling et al. (2022) and Huang et al. (2023b), we evaluate SIDSL on four real-world propagation datasets containing actual cascades: Digg, Twitter, Android, and Christianity. For each cascade, we define source nodes as those infected within the first 10% of the propagation time and use the network’s infection status at 30% of the time as the observation input. This setup simulates real-world applications, where initial events must be inferred from data collected after a delay. Further dataset details are in Appendix A.
Implementation Details. We use a 6:1:1 train/validation/test split for all datasets. The diffusion framework employs $T=500$ timesteps with linear noise scheduling. The denoising network uses a 2-layer GCN for LPSI estimation $X_{est}$ conditioning, while the GNN encoder/decoder comprises 3-layer GCNs (hidden dimension 128) and the residual GNN uses 2 layers (hidden dimension 8). We use Adam optimizer with learning rate searched from 0.01, 0.005, 0.001, linear decay scheduling, and maximum 500 training epochs.
Baselines. Following previous works (Ling et al., 2022; Yan et al., 2024), we selected two representative heuristic methods, i.e., Netsleuth (Prakash et al., 2012) and LPSI (Wang et al., 2017), and deep learning methods, i.e., GCNSI (Dong et al., 2019), TGASI (Hou et al., 2023) and recent generative deep learning methods SLVAE (Ling et al., 2022), DDMSL (Yan et al., 2024) that capture source localization uncertainty. These baselines are all state-of-the-art (SOTA) multi-source localization methods in their domains. Please refer to the Appendix C for specific information on baselines and our method.
Metrics. Following (Wang et al., 2022), we use three metrics: 1) F1-score (F1): The harmonic mean of precision and recall; 2) Recall (RE): The proportion of correctly identified positive cases (source nodes); and 3) Precision (PR): The proportion of true positives among predicted positive cases. We omit accuracy because it is a misleading metric in highly imbalanced datasets, as it fails to capture the model’s ability to identify the rare positive class.
@@ 5.2. Overall Performance on Real-world Datasets
To evaluate the real-world performance of our proposed method against baselines, we conduct direct training and testing on four real-world datasets for all methods. The experimental results are presented in Table 1. Our proposed SIDSL outperforms all baselines across nearly all metrics on all datasets. Specifically, SIDSL achieves F1 scores that exceed the second-best baseline by 7.5%, 11.0%, 11.8%, and 13.3% on Digg, Twitter, Android, and Christianity, respectively, which demonstrates SIDSL’s superior ability to predict source nodes despite their sparsity. Our model also excels in recall, with significant gains in Android (48.3%) and Christianity (12.8%), demonstrating strong ability in identifying true source nodes.
Further, we have the following findings. First, compared to purely data-driven generative methods like DDMSL and SLVAE, SIDSL demonstrates superior performance, validating the effectiveness of incorporating structural prior information. Second, compared to rule-based and learning-based methods, generative methods (SIDSL, SLVAE, DDMSL) show stronger performance, highlighting the advantages of effective source distribution modeling. LPSI and other deep learning-based methods perform comparably well, successfully capturing structural characteristics. Netsleuth demonstrates the weakest performance due to limited generalization beyond SI propagation patterns (Keeling and Eames, 2005; Prakash et al., 2012). Finally, SIDSL shows the most substantial improvement on the Christianity dataset (13.3% in F1), as a smaller network scale enhances topological feature capture, amplifying structural prior benefits.
@@ 5.3. Performance in Low Data Regime
To assess how deep learning methods leverage simulation data pretraining for few-shot and zero-shot learning in real-world source localization, we generate pretraining data using standard IC and LT propagation models (1:1 ratio) across four networks. These models effectively capture social network dynamics by representing real-world interaction randomness, cumulative influence, and topology-based spread patterns. For few-shot learning (P), we randomly selected 10% of the training data, sharing this data across all evaluated methods. We also compare the performance of all the methods trained on equivalent data volumes but without pretraining (NP). For zero-shot learning, all the methods are directly tested after being pretrained.
@@
Few-shot analysis. The result of few-shot learning is shown in Tabel 2. Rule-based methods (Netsleuth and LPSI) are omitted as they are training-free. SIDSL demonstrates superior performance compared to all baseline methods across all datasets, both with and without pretraining, achieving improvements of 24-365% across all metrics. We observe two key findings. First, SIDSL demonstrates enhanced performance advantages when leveraging pretraining, with 8 out of 12 cases showing larger gains with P compared to NP across different datasets and metrics. This pattern of larger improvements with pretraining demonstrates SIDSL’s superior ability to effectively leverage information from simulated data to real-world scenarios in limited-data settings. Second, baseline methods exhibit varying capabilities in utilizing pretraining. While TGASI shows consistent improvements with pretraining, generative approaches (SLVAE, DDMSL) demonstrate limited gains (e.g., F1 increases from 0.009 to 0.036 for SLVAE, 0.010 to 0.109 for DDMSL on Android). This highlights generative models’ sensitivity to distribution shifts, whereas our method achieves robust improvements through effective incorporation of pattern-invariant structural priors.
@@
Zero-shot analysis. The results of zero-shot learning are shown in Table 3. SIDSL significantly outperforms all baseline methods across all datasets, achieving improvements up to 364% in F1 scores. SIDSL achieves particularly strong performance on Android and Christianity datasets (¿0.9 recall with reasonable precision). Baseline methods show limited zero-shot capability (F1 ¡ 0.3). Zero-shot SIDSL outperforms rule-based methods (Table 1) in most metrics. These findings demonstrate SIDSL’s superior simulation-to-real transfer without fine-tuning.
In few-shot and zero-shot settings, pretrained SIDSL achieves higher recall than precision gains, effectively identifying true sources but with occasional false positives, especially in Android and Christianity datasets. This suggests strong propagation modeling despite some errors. For misinformation tracing, pretrained SIDSL provides superior source identification. Accuracy can be enhanced using precision-based classifiers with boosting to filter false positives. Without pretraining, SIDSL outperforms other methods on both metrics, proving effective in low-data scenarios.
@@
Figure 4. Few-shot performance of SIDSL with different pretraining dataset compositions (IC+LT, IC, LT), then finetuned on real-world data. ”NP” denotes ”non-pretrained”.